OpenStudy (x__hazel__x):

What two numbers multiply together to give you -72 but add to give you 10

OpenStudy (anonymous):

\[ ab=-72\\ a+b=10 \]

OpenStudy (x__hazel__x):

Thank you!

OpenStudy (anonymous):

This actually is very much like the quadratic equation:\[ (x+a)(x+b) = x^2+(a+b)x+ab = x^2+10x-72 \]

OpenStudy (x__hazel__x):

wowoo! Thank youuu!

OpenStudy (anonymous):

Do you know how to solve a quadratic equation?

OpenStudy (x__hazel__x):

yes, thank you!